# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ VI SOTA in Direct AI/ML Positioning
  • Channel charting based positioning: Channel charting based positioning uses self-supervised learning to model relationships between signal sampling points, creating a pseudo-coordinate system from raw CSI data without relying on absolute positions or detailed environment knowledge. It is scalable and reduces dependency on labeled datasets, making it suitable for dynamic and large-scale scenarios.
@@ VI-A1 Fingerprint Collection and Feature Extraction
In the fingerprint collection and feature extraction stage, wireless signals are collected offline at RPs, and relevant features are extracted to serve as fingerprints. The simplest forms of fingerprints include RSS [258, 259, 260, 261, 262, 263], RSRP [264], RSRQ [265, 264], and TOA measurements [145, 146]. However, these types of fingerprints are often insufficiently detailed and lack the reliability required for positioning performance due to factors such as device errors and measurement noise.
CSI, with its higher-dimensional signal representation, captures the essential characteristics of wireless signal transmission, encapsulating comprehensive information about the wireless environment. This makes CSI a valuable resource for achieving high-precision positioning. However, the high-dimensional nature of CSI also poses challenges. First, extracting location-related features from raw CSI data is not easy because it is usually difficult to define which features are more important for the positioning task. Second, higher-dimensional CSI also increases the complexity of AI algorithms.
Therefore, signal preprocessing of CSI is crucial to effectively extract meaningful positioning features from raw CSI data. Some works directly utilize CFR [266, 267, 268] as model input. To extract features like AOA and TOA more effectively, several studies apply the discrete fourier transform (DFT) [269, 270, 271, 272] for CIR feature extraction. Furthermore, techniques such as quantization, normalization, grayscale processing, and principal component analysis [269] are employed to reduce parameter count and save storage space. For example, in [273], the paper combines RSS and CSI, using CNN-based feature fusion and gradient blending to improve accuracy. In [274], the paper proposes a positioning framework based on deep learning under commercial LTE systems, utilizing real-time CFR data collected via software-defined UE and employing a time-domain fusion approach to enhance positioning accuracy and robustness. The authors in [275] extract the energy coupling matrix in the refined beam domain as CSI-based fingerprints. In [271], DFT-extracted angle-domain channel power matrices are used as fingerprints, with spectral clustering for efficient matching and weighted K-nearest neighbors (WKNN) for location estimation. The authors in [276] propose a CSI quality control module that combines the Hampel identifier, wavelet filter, and cross-correlation detector to obtain clean and stable 5G NR CSI fingerprints, as well as a linear transformation module to mitigate the effects of sampling frequency offset and carrier frequency offset. In [277], the authors utilize an angle-delay channel power matrix (ADCPM) as a high-resolution fingerprint, complemented by a novel similarity criterion, compression, and two-stage clustering for efficient database preprocessing and improved localization performance. Similarly, the paper in [95] also uses the sparsity-enhanced ADCPM fingerprints to extract multipath characteristics and improve positioning accuracy while reducing computational complexity and noise sensitivity.
[…]
@@ VI-A2 Fingerprint Database Construction
Traditional fingerprint positioning relies on constructing a fingerprint database and determining user locations by comparing real-time measurements with the database. Interpolation methods are widely used across time, frequency, and spatial domains to reconstruct complete radio maps for positioning [279, 280]. Model-based interpolation techniques, such as linear interpolation [281], Gaussian process regression [282], Kriging interpolation [283], and Voronoi tessellation-based interpolation method [284] have been extensively utilized for this purpose. For example, in [281], the paper explores approximating RSSI values using linear interpolation and Gaussian process regression, balancing accuracy, computational complexity, and data collection time. Advanced approaches also include multicomponent optimization and sparse recovery [285], total variation norm minimization for edge-preserving ray tracing [286], and energy-modified leverage sampling to improve matrix completion [287]. To further enhance the performance of the radio map, other methods leverage positional uncertainty [282], propagation priority [288], and graph-based signal processing incorporating NLOS conditions [289]. All of these techniques attempt to construct reliable and adaptable radio maps from the perspectives of accuracy, complexity, and data efficiency.
In addition, AI-based interpolation methods have also been widely studied. In [290], the paper proposes a novel deep learning framework for radio map construction, transforming spatial interpolation into a shadowing adjustment problem and introducing a neural network structure suitable for this task. In [291], the authors propose an LSTM-based deep learning method for radio environment map reconstruction in V2X scenarios. The authors in [292] propose a super-resolution radio environment map construction method combining Kriging interpolation, dictionary learning, and Random Forest, enabling efficient and accurate high-resolution radio environment maps with reduced computation time. Similarly, the study in[293] investigates feedforward neural networks for path loss modeling to enhance Kriging-based radio environment mapping. The authors [294] propose a two-phase learning framework integrating radio propagation models with a conditional GAN to extract global propagation patterns and local shadowing effects. Additionally, DeepREM uses U-Net and conditional GAN models to estimate the radio environment map from sparse measurements without requiring any additional information [295]. Moreover, in [296], the paper leverages a fully convolutional deep completion autoencoder to learn spatial propagation structures from prior data, significantly reducing measurement requirements while enhancing accuracy. LocUNet [297] uses an end-to-end CNN for user localization using RSS from a small number of BSs, leveraging pathloss radio map estimations for robust real-time performance without the need for specific area fingerprints.
@@
However, the fingerprint positioning approach based on the radio environment map depends heavily on the density of RPs and the efficiency of the query algorithm. It faces challenges such as high deployment costs and prolonged query times. With the advancement of AI technologies, existing studies increasingly replace fingerprint databases with AI models. AI algorithms learn the mapping between wireless fingerprints and locations from datasets, significantly improving positioning accuracy. This enhancement in accuracy largely depends on the design of the AI models, leading to the development of effective models using architectures like FCNN [265], CNN [261, 266, 298, 95], LSTM [262], RNN [299, 300], and Transformers [301, 269]. For example, in [302], the paper proposes a hybrid indoor positioning architecture combining CNN, LSTM, and GAN models to enhance training data and improve accuracy. DeepWiPos uses an LSTM-based framework with attention modules to fuse RSS and fingerprint spatial gradients, addressing RSS instability and spatial ambiguity [303]. The paper in [298] investigates Residual Block-based CNN algorithms, using uplink beamformed CSI fingerprints with multiple spatial dimensions at both the BS and UE. Furthermore, in [301], the authors propose CrowdBERT, a transformer-based semi-supervised algorithm that leverages crowdsourced fingerprint data and employs spatial attention encoding, RSS-token masking, and fine-tuning to enhance spatial feature extraction.
To further enhance the robustness of the algorithm, some studies leverage advanced AI techniques such as contrastive learning [304], semi-supervised learning [305, 306], autoencoders[276], and meta-learning-based [307, 308] to improve positioning accuracy. In [304], the authors employ contrastive learning to capture wireless channel representations. By applying stochastic channel augmentations, they generate different views of the channel and learn representations for both micro-fading and macro-fading effects. In [305], a semi-supervised learning framework using a signal-guided masked autoencoder is proposed for high-precision positioning with limited samples. Additionally, an LSTM network is incorporated to leverage CIR’s temporal characteristics for user coordinate estimation. In [306], the authors propose a semi-supervised contrastive learning technique for massive MIMO fingerprint positioning, leveraging partially labeled pilot signal data and data augmentation to pre-train an encoder with contrastive loss, followed by fine-tuning for accurate positioning. The authors in [276] propose the iPos5G system, which first uses an unsupervised deep autoencoder network to reconstruct CSI features, then employs supervised learning to improve the RBF for optimizing the similarity calculation probability model, and finally performs positioning using an amplitude-phase probability fusion function. In [307], the paper proposes a Bayesian meta-learning-based fast adaptation approach to address the challenge of outdated samples, enabling pre-trained models to quickly adapt to new tasks with improved robustness. The authors in [264] propose 5G1M, a simplified 5G-based indoor positioning algorithm leveraging a Siamese network with Ghost modules, transfer learning, and trajectory fitting to reduce database dependence and enhance adaptability. Furthermore, in [309], the paper compares early and late fusion techniques, proposing a multi-task learning scheme to improve accuracy and efficiency while leveraging uncertainty estimation for enhanced fusion reliability.
To further reduce the cost of model construction, transfer learning and similar methods are utilized to migrate existing models and knowledge from one environment to another, thereby lowering deployment overhead. In [267], the paper proposes a meta-learning-based deep learning model for radio-based positioning, enabling improved transfer learning by separating environment-independent feature learning from environment-specific adaptation. TransLoc [41] is a heterogeneous knowledge transfer framework for fingerprint-based indoor localization. It addresses real-time environmental dynamics by refining source domain knowledge, creating a homogeneous feature space through cross-domain mapping, and utilizing a joint optimization algorithm for efficient knowledge transfer with minimal target domain samples. In [310], the study addresses cross-environment indoor positioning using a semi-supervised approach, introducing a deep neural forest combined with adversarial training to learn environment-invariant features, enabling robust localization without requiring annotations in new environments. Moreover, the authors in [272] present the cross-region fusion and fast adaptation framework for fingerprint localization in cell-free massive MIMO systems, leveraging cross-region fusion, meta-learning, and fine-tuning to enhance localization accuracy, achieve rapid deployment.
@@ VI-A3 Fingerprint Database Updating
Fingerprint database and AI model aging also pose critical challenges in wireless fingerprint-based positioning. Environmental changes, such as infrastructure updates, furniture rearrangements, or time variations, can cause the database and models to become outdated, reducing positioning accuracy. Addressing this issue requires efficient methods for maintaining and updating the database and AI models.
Crowdsourcing methods provide an effective solution by opportunistically collecting labeled fingerprints during routine device usage. Crowdsourced fingerprint updates leverage opportunistic real-world data, such as high-confidence labels obtained near base stations or in areas with strong GPS signals [311, 312, 313], as well as user movement trajectories obtained from pedestrian dead reckoning (PDR) [314, 315, 316, 317, 318, 142, 319]. For example, the authors in [311] use multi-kernel transfer learning for integrating historical fingerprints and GPS-based opportunistic fingerprints to achieve one-meter accuracy with minimal update overhead. Based on trajectory continuity, the authors in [315] propose integrating a modified particle filter with a self-updating fingerprint database to enhance accuracy and robustness. Similarly, in [317], the paper leverages static mobile devices as RPs and a trajectory-matching algorithm to adapt the radio map to dynamic environments. In addition, the authors in [320] propose a crowdsourcing-based radio map update method using sparse representation and low-rank matrix recovery to address incomplete and noisy fingerprints. Moreover, managing computational demands and filtering crowdsourced data also pose significant challenges. In [321], the paper transforms crowdsourcing trace data into cluster and positioning spaces, enabling unsupervised online updates with dynamic replay memory. However, these crowdsourcing-based methods all rely on additional data such as GPS and PDR, and it is still challenging to accurately determine the location of each crowdsourced fingerprint without prior information.
Advanced techniques like continual learning [48, 322] and GANs [323] have also played a pivotal role in addressing database aging. In [323], the authors introduce RecTrack-GAN, an indoor tracking framework combining RNNs for human movement modeling and Conditional GANs for data recovery and fingerprint database updates, improving tracking accuracy. In [48], the authors propose an incremental learning-based fingerprint localization scheme using CSI images and a broad learning system, enabling rapid updates without retraining and achieving superior performance in real-world indoor environments. In [324], the paper proposes a continuous active learning method and uses uncertainty sampling to periodically adapt the database to environmental changes, improving accuracy and preventing performance degradation. In addition, based on transfer learning, the authors in [322] propose a system that automatically updates radio maps using transfer learning with altered AP identification and mapping space searching, significantly improving accuracy in dynamic environments. The authors in [308] present FeMLoc, a federated meta-learning framework for indoor localization, enabling swift adaptation to new environments with minimal data and reduced calibration effort. The authors in [268] introduce a semantic localization approach enhanced by multi-task deep domain adaptation and scenario adaptive learning, integrating environmental semantics into localization frameworks to address time-varying signal propagation challenges.
@@ VI-B Knowledge-assisted AI Positioning
Traditional positioning methods, such as trilateration, triangulation, and multilateration, utilize geometric principles and channel propagation models to estimate position based on signal properties. While these geometry-based methods are effective under some conditions, they often face significant challenges in complex or dynamic environments, particularly in scenarios affected by multipath effects and NLOS conditions.
In contrast, fingerprint-based positioning uses AI to directly learn the mapping between signal characteristics (or features) and the position, achieving end-to-end positioning. However, as noted above, fingerprint positioning systems can experience performance degradation due to changes in the wireless environment. These systems also face significant difficulties in transferring across different environments or deployment scenarios, which limits their robustness and scalability.
To address the limitations of both traditional geometry-based and pure fingerprint-based positioning approaches, knowledge-assisted AI positioning has emerged. This approach combines the power of AI with domain knowledge, such as geometric and channel propagation insights, enabling positioning systems to achieve greater accuracy and reliability. Unlike purely data-driven, end-to-end methods like fingerprinting, knowledge-driven AI-assisted positioning integrates physical models and environment-specific features as a foundation for AI models. This fusion of knowledge allows AI algorithms to perform effectively, even in challenging environments where data alone may fall short.
Through this combination of AI with domain knowledge, knowledge-assisted AI positioning can adapt more robustly to environmental variability and reduce reliance on extensive, labeled datasets, which are often required for purely data-driven approaches. This integration enhances accuracy and improves generalization across various environments, making the positioning system more resilient to external changes. In this subsection, we will introduce the research related to wireless-knowledge-assisted positioning and geometric-knowledge-assisted positioning respectively.
@@ VI-B1 Wireless-knowledge-assisted Positioning
A considerable amount of research focuses on integrating wireless domain knowledge into AI algorithms to optimize positioning systems. Combined with this wireless domain knowledge, AI algorithms can effectively solve problems such as wireless channel feature extraction [325, 326, 327] and hardware damage [328, 329], thereby improving algorithm robustness. Several studies exploit the intrinsic properties of wireless channels to enhance the robustness of AI-based positioning [325, 326, 327, 304, 330, 331]. For instance, to address frequency diversity challenges in wireless networks, the authors in [325] propose a multi-frequency fusion learning method. This approach first extracts position-related features independently from CSI at each subcarrier using a shareable method and then fuses these features to overcome the diversity challenges inherent to FDMA systems. In [326], a joint approach considers both frequency-domain and time-domain features. In the frequency domain, relative phase differences and received powers across resource blocks are combined, while in the time domain, parameters like AOA, RTT, and signal strength are incorporated to jointly improve positioning accuracy. Recognizing the high dimensionality of channel data, the authors in [327] propose a positioning neural network and reduce computational complexity by utilizing minimal descriptive features, including maximum power measurements and their temporal locations. In [330], the authors present a view-selective deep learning system, leveraging multiview training with supervised variational networks and dominant view classification to enhance localization accuracy in complex radio environments. Additionally, in distributed M-MIMO frameworks, the study [331] introduces a deep belief network to analyze RSS values derived from diffraction models for positioning.
@@
Considering hardware impairments, the authors in [328] explore end-to-end positioning using an autoencoder architecture, effectively mitigating hardware-induced errors. In [329], the authors address challenges specific to RIS-aided mmWave systems, such as clock offsets between transmitters and receivers, impairments at the transmit and receive antenna arrays, and coupling effects among RIS elements. They propose a dictionary learning approach to calibrate these hardware impairments, significantly improving positioning performance in RIS-aided environments. These methods demonstrate the importance of accounting for hardware imperfections in achieving accurate and robust localization.
In addition, based on communication system properties, some studies explore the synergies between wireless positioning and other communication tasks, such as channel estimation [329] and CSI feedback [332]. For instance, the authors in [329] jointly optimize channel estimation and localization in RIS-aided mmWave systems. In [332], an integrated learning framework for CSI feedback and localization is proposed. This framework enables both tasks to mutually benefit each other, demonstrating that incorporating coarse positional data improves the accuracy of both CSI feedback and CSI-based localization.

### v2

@@ VI SOTA in Direct AI/ML Positioning
  • Channel charting based positioning: Channel charting based positioning uses learning methods to model relationships between signal sampling points, creating a pseudo-coordinate system from raw CSI data without relying on absolute positions or detailed environment knowledge. It is scalable and reduces dependency on labeled datasets, making it suitable for dynamic and large-scale scenarios.
@@ VI-A1 Fingerprint Collection and Feature Extraction
TABLE X: Overview of Signal Types and Representations Used as Fingerprints in Localization
Signal / Representation Representative Works
RSS / RSRP / RSRQ [145, 269, 270, 271, 272, 273, 274, 275]
TOA / AOA Measurements [109, 110]
CFR / Spatial-Frequency Domain CSI [276, 277, 278]
CIR / Angle-Delay Domain CSI [279, 280, 281, 170, 138, 61, 282]
Filtered / Preprocessed CSI [283, 284]
Fingerprint Fusion (e.g., RSS, CSI) [285, 286]
Sensor Fusion (e.g., WiFi, IMU, GNSS) [287, 272, 273]
In the fingerprint collection and feature extraction stage, wireless signals are collected offline at RPs, and relevant features are extracted to serve as fingerprints. The simplest forms of fingerprints include RSS [145, 269, 270, 271, 272, 273], RSRP [274], RSRQ [275, 274], and TOA/AOA measurements [109, 110, 288]. However, these types of fingerprints are often insufficiently detailed and lack the reliability required for positioning performance due to factors such as device errors and measurement noise.
CSI, as a high-dimensional representation of wireless channels, provides rich spatial and temporal information for positioning. Various forms of CSI-based fingerprints have been explored in the literature. Some studies directly use the CFR as input, with either amplitude–phase or real–imaginary components encoded as two separate feature maps [276, 277, 278]. To better capture high-resolution multipath characteristics, the discrete Fourier transform (DFT) is often applied to obtain CIR or angle-delay domain features such as the ADCPM [170, 279, 280, 281, 138]. Building upon this, technologies such as sparsity-enhanced ADCPM and refined beam domain channel matrices have been proposed to improve multipath discrimination and robustness [61, 282]. Given the presence of noise and hardware imperfections in practical systems, several works also explore preprocessing techniques for CSI to enhance the learnability and stability of fingerprints [283, 284]. Moreover, recent studies demonstrate that fusing different types of wireless fingerprints (such as CSI and RSS) [285, 286], or integrating auxiliary sensors (like IMUs and GNSS) [287, 272, 273] can further improve positioning accuracy and robustness, particularly in complex environments. Table. X summarizes representative fingerprint types used in wireless positioning.
@@ VI-A2 Fingerprint Database Construction
TABLE XI: Overview of Fingerprint Database Construction Methods
Method Category Representative Works
Model-based interpolation method for radio map construction Linear interpolation [289], Kriging interpolation [290, 291, 292], Gaussian process regression [293], Voronoi tessellation based interpolation[136] and multicomponent optimization and sparse recovery [294].
AI-assisted interpolation method for radio map construction FCNN [295], CNN [296], LSTM [297], Transformer [298], GAN [299, 143], and autoencoder [300]
AI-driven positioning based on fingerprint mapping FCNN [275], CNN [271, 276, 301, 61], LSTM [272], RNN [302, 303], and Transformers [304, 170].
@@
An overview of fingerprint database construction methods is provided in Table. XI. Traditional fingerprint positioning relies on constructing a fingerprint database and determining user locations by comparing real-time measurements with the database. Interpolation methods are widely used across time, frequency, and spatial domains to reconstruct complete radio maps for positioning [290, 291]. Model-based interpolation techniques, such as linear interpolation [289], Kriging interpolation [290, 291, 292], gaussian process regression [293], Voronoi tessellation-based interpolation[136], as well as multicomponent optimization and sparse recovery [294] have been extensively utilized for this purpose. All of these techniques attempt to construct reliable and adaptable radio maps from the perspectives of accuracy, complexity, and data efficiency. In parallel, AI-based interpolation methods have attracted increasing attention. These methods leverage a variety of learning models, including FCNNs [295], CNNs [296], LSTMs [297], and Transformers [298]. In addition, GANs [299, 143] are employed to synthesize fingerprint data for unmeasured positions through adversarial learning, while autoencoders [300] compress sparse radio maps into low-dimensional latent spaces and reconstruct interpolated results through decoding. These AI-driven approaches offer greater modeling flexibility and are better suited to capturing complex, nonlinear spatial variations in wireless signal propagation.
@@
However, the fingerprint positioning approach based on the radio environment map depends heavily on the density of RPs and the efficiency of the query algorithm. It faces challenges such as high deployment costs and prolonged query times. With the advancement of AI technologies, existing studies increasingly replace fingerprint databases with AI models. AI algorithms learn the mapping between wireless fingerprints and locations from datasets, significantly improving positioning accuracy. This enhancement in accuracy largely depends on the design of the AI models, leading to the development of effective models using architectures like FCNN [275], CNN [271, 276, 301, 61], LSTM [272], RNN [302, 303], and Transformers [304, 170]. For example, in [305], the paper proposes a hybrid indoor positioning architecture combining CNN, LSTM, and GAN models to enhance training data and improve accuracy. DeepWiPos uses an LSTM-based framework with attention modules to fuse RSS and fingerprint spatial gradients, addressing RSS instability and spatial ambiguity [306]. The paper in [301] investigates Residual Block-based CNN algorithms, using uplink beamformed CSI fingerprints with multiple spatial dimensions at both the BS and UE.
TABLE XII: Summary of AI-driven Fingerprint Update Works
Update Type Ref. Algorithm Label Needed Time/Space Shift Key Insight
Radio Map Update [307] GAN Yes Time Use conditional GAN to generate time-aware RSSI maps for updating fingerprints.
[308] Few-shot Learning + GAN Yes Time Augment fingerprints via GAN and infers location labels by spectral clustering.
[309] Incremental Learning No Time Use incremental learning adapt to time-varying CSI without retraining CNN.
[310] Model-based Transfer Learning Yes Time Align heterogeneous RSS/CSI domains via feature mapping and weighted instance transfer without retraining.
[142] Model-based Transfer Learning Yes Time Use enhanced transfer component analysis to align domains and filters changed APs for radio map updating.
Neural Network Update [311] Transfer Learning Yes Time + Space Handle label sparsity and domain shifts via Bayesian domain transfer for cross-domain positioning.
[312] Transfer Learning + Knowledge Distillation Yes Time + Space Combine knowledge distillation with fine-tuning on real data to adapt from synthetic to real scenarios.
[313] Transfer Learning Yes Time + Space Use dual-domain adaptation and weight regularization to adapt CSI-based model across dynamic conditions.
[281] Meta-Learning Yes Space Use cross-region fusion and meta-Learning to enable model adaptation.
[176] Meta-Learning Yes Space Learn environment-agnostic meta-parameters for rapid adaptation with few labeled samples.
[314] Federated Meta-Learning Yes Time + Space Learn a meta-model via federated training to rapidly adapt to new environments.
[177] Bayesian Meta-Learning Yes Time Combine Bayesian meta-learning to enable fast adaptation of localization.
@@ VI-A3 Fingerprint Database Updating
Fingerprint database and AI model aging also pose critical challenges in wireless fingerprint-based positioning. Environmental changes, such as infrastructure updates, furniture rearrangements, or time variations, can cause the database and models to become outdated, reducing positioning accuracy. Addressing this issue requires efficient methods for maintaining and updating the database or AI models. To achieve this goal, crowdsourcing methods provide an effective solution by opportunistically collecting labeled fingerprints during routine device usage. Crowdsourced fingerprint updates leverage opportunistic real-world data, such as high-confidence labels obtained near base stations or in areas with strong GPS signals [141, 315, 316], as well as user movement trajectories obtained from pedestrian dead reckoning (PDR) [317, 318, 319, 320, 321, 106, 322]. With the help of crowdsourced data, we have the opportunity to update the fingerprint to maintain positioning accuracy.
Advanced AI techniques have been widely applied to tackle the challenges of radio map aging and domain adaptation in fingerprint-based localization. Fingerprint databases built on RSS or CSI measurements may degrade over time or vary across environments. To address this, several works employ radio map updates using GANs to synthesize new fingerprints [307, 308] or transfer learning to adapt old fingerprints to new domains [310, 142, 311]. In [309], incremental learning has also been used to continuously adapt fingerprints without retraining. In contrast, when deep models replace explicit fingerprint databases, model adaptation becomes essential. Methods based on transfer learning [312, 313], meta-learning [281, 176, 177], and federated learning [314] aim to fine-tune neural networks with minimal labeled data. The comparison of related work is shown in Table. XII.
@@ VI-B Knowledge-assisted AI Positioning
Model-based positioning methods, such as trilateration, triangulation, and multilateration, utilize geometric principles and channel propagation models to estimate position based on signal properties. While these geometry-based methods are effective under some conditions, they often face significant challenges in complex or dynamic environments, particularly in scenarios affected by multipath effects and NLOS conditions. In contrast, fingerprint-based positioning uses AI to directly learn the mapping between channel (or features) and the position, achieving end-to-end positioning. However, as noted above, fingerprint positioning systems can experience performance degradation due to changes in the wireless environment. These systems also face significant difficulties in transferring across different environments or deployment scenarios, which limits their robustness and scalability.
To overcome these challenges, knowledge-assisted AI positioning integrates domain knowledge, such as geometry and channel models, into AI-driven approaches. By embedding physical insights into learning models, this hybrid method improves positioning accuracy and generalization, even in variable environments, while reducing the dependency on large labeled datasets. This subsection introduces two key branches of this approach: wireless-knowledge-assisted positioning and geometric-knowledge-assisted positioning.
@@ VI-B1 Wireless-knowledge-assisted Positioning
TABLE XIII: Categorization of Wireless-Knowledge-Assisted Positioning Methods
Category Methods References
Knowledge from wireless system characteristics Knowledge derived from network configurations [163, 323]
Knowledge derived from different channel-related tasks [135, 324]
Knowledge from unlabeled wireless data Generative-based SSL pretraining [325, 304, 201]
Contrastive-based SSL pretraining [201, 326, 327, 328, 169]
Knowledge-driven mitigation for deployment imperfections Deployment imperfections such as hardware impairments, calibration errors, and clock offsets. [329, 135]
Wireless-knowledge-assisted positioning aims to enhance AI-based localization by embedding domain knowledge of wireless propagation into learning models. To systematically summarize existing approaches, Table. XIII categorizes representative methods based on the type of wireless knowledge utilized and its functional role in the positioning process. One effective strategy involves designing customized localization algorithms by leveraging the system characteristics of wireless networks. On one hand, knowledge derived from specific network configurations can be exploited to improve positioning accuracy [163, 323]. For instance, in FDMA systems, multi-frequency fusion has been proposed to harness subcarrier diversity for more robust localization [163]. On the other hand, system-level knowledge can be utilized through joint design with other communication tasks, such as channel estimation [135] and CSI feedback [324], which share underlying characteristics with localization. For example, [135] jointly optimizes channel estimation and localization in RIS-aided mmWave systems. Similarly, [324] proposes an integrated learning framework in which CSI feedback and localization mutually benefit from each other.
Another way to incorporate wireless knowledge is by learning channel representations from large volumes of unlabeled data using self-supervised learning (SSL). SSL enables models to capture the intrinsic semantic structure of the wireless channel, such as spatial geometry, multipath propagation patterns, and user mobility characteristics, without relying on costly location labels. SSL approaches can be broadly categorized into two paradigms: generative-based methods (e.g., masked modeling) [325, 304, 201], and contrastive-based methods [201, 326, 327, 328, 169]. Once pre-trained using these strategies, the learned channel representations can be fine-tuned with a small amount of labeled data to achieve significantly improved localization performance [169, 201]. In particular, the authors in [169] propose a Large Wireless Localization Model (LWLM) that combines generative-based and contrastive-based SSL objectives. By pretraining on large-scale CSI data, the model achieves performance gains across various localization tasks, demonstrating the effectiveness in capturing channel semantics relevant to positioning.
@@
Beyond improving localization accuracy, wireless domain knowledge is also leveraged to address practical challenges in real-world wireless systems, such as hardware impairments, calibration errors, and clock offsets, all of which can significantly degrade positioning performance during deployment [329, 135]. Considering hardware impairments, the authors in [329] explore end-to-end positioning using an autoencoder architecture, effectively mitigating hardware-induced errors. In [135], the authors address challenges specific to RIS-aided mmWave systems, such as clock offsets between transmitters and receivers, impairments at the transmit and receive antenna arrays, and coupling effects among RIS elements.